### v1

# From Estimands to Robust Inference of Treatment Effects in Platform Trials

Abstract
A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner and can accelerate the evaluation of new treatments. However, the flexibility that marks the potential of platform trials also creates inferential challenges. Two key challenges are the precise definition of treatment effects and the robust and efficient inference on these effects. To address these challenges, we first define a clinically meaningful estimand that characterizes the treatment effect as a function of the expected outcomes under two given treatments among concurrently eligible patients. Then, we develop weighting and post-stratification methods for estimation of treatment effects with minimal assumptions. To fully leverage the efficiency potential of data from concurrently eligible patients, we also consider a model-assisted approach for baseline covariate adjustment to gain efficiency while maintaining robustness against model misspecification. We derive and compare asymptotic distributions of proposed estimators in theory and propose robust variance estimators. The proposed estimators are empirically evaluated in a simulation study and illustrated using the SIMPLIFY trial. Our methods are implemented in the R package RobinCID.
Keywords: Covariate adjustment; Eligible and concurrent controls; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
1 Introduction
The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Saville and Berry,, 2016; Woodcock and LaVange,, 2017). As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
The platform trial has emerged as an innovative trial design to expedite the evaluation of potentially effective treatments (Berry et al.,, 2015; Saville and Berry,, 2016; Woodcock and LaVange,, 2017; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023). It is a dynamic clinical trial that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner, where new treatments can be added to the platform as they become available, and existing ones can leave the platform based on a finding of futility or efficacy. Such a trial structure not only streamlines operations but also has the potential to improve statistical efficiency by facilitating the sharing of data between different treatment evaluations (Mehta et al.,, 2023). The effectiveness of platform trials in addressing urgent medical challenges quickly is evidenced by the hallmark RECOVERY platform trial, which was designed to rapidly evaluate COVID-19 therapies during the pandemic (Normand,, 2021; Buenconsejo et al.,, 2023). Platform trials are also attractive in the exploratory phase to quickly screen multiple treatments for diseases (Wang and Yee,, 2019) and in rare diseases to enhance the chance of identifying effective treatments (Dhaenens et al.,, 2024). Two examples of platform trials are presented in Section 2 to provide detailed illustrations.
Although platform trials offer several advantages over traditional trials, they also create challenges in statistical inference on treatment effects, as highlighted in a recent Food and Drug Administration draft guidance (FDA,, 2023). One primary challenge is how to define a meaningful treatment effect in a platform trial setting, as multiple treatments may target different biomarker-defined subgroups and enter the trial at different times — a complexity that does not exist in traditional trials. In platform trials, when comparing a control and a treatment of interest, many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, (2023) have recommended using eligible and concurrent controls, which refer to the patients who are eligible to receive the treatment and are concurrently randomized to the control arm during the same enrollment period as the other patients being assigned to the treatment arm. The use of eligible and concurrent controls, or concurrently eligible patients, preserves the integrity of randomized comparisons and inference on treatment effects. In contrast, patients who are either ineligible or nonconcurrent may not be comparable to the treated group due to measured and unmeasured differences (FDA,, 2023). In Section 3, we explicitly present, for the first time in the literature, how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients, which aligns with the ICH E9 , 2019 (R1) estimand framework.
The next challenge is on robust estimation of the defined treatment effects for concurrently eligible patients. Given the adaptive nature of platform trials, where treatments can be added and removed, the available treatment options vary over time. This dynamic, coupled with different eligibility criteria for each treatment, means that the probabilities of treatment assignments can vary over time and patient characteristics — another complexity absent in traditional trials. Thus, naively comparing outcome means between two treatment arms may yield biased results, an example of Simpson’s paradox (Collignon et al.,, 2020; Park et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021).
The literature on statistical methods for platform trials is expanding, though it primarily focuses on model-based techniques for incorporating nonconcurrent controls, which can be sensitive to model misspecification (Saville et al.,, 2022; Bofill Roig et al.,, 2022; Elm et al.,, 2012; Wang et al.,, 2023). Some research has been conducted on statistical methods for concurrently eligible patients. Ventz et al., (2018) considered several adaptive randomization procedures and tests for the hypothesis of no treatment effect. Marschner and Schou, (2022) proposed stratification using enrollment windows, without taking into account other factors related to treatment and randomization. Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows. However, there remains a lack of rigorous investigation on how to make robust inference on treatment effects in platform trials under the same minimal statistical assumptions as those in traditional clinical trials.
To fill this gap, in Section 4, we develop methods for estimating treatment effects for concurrently eligible patients using inverse probability weighting or post-stratification to account for varying treatment assignment probabilities. Additionally, we consider model-assisted approaches for covariate adjustment to gain efficiency while maintaining robustness against model misspecification. In Section 5, we theoretically establish the asymptotic distributions of proposed estimators and compare their asymptotic relative efficiency. We provide robust variance estimators in Section 6. Simulation results and an application of our methods to the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) are in Section 7. Our methods are implemented in the R package RobinCID.
2 Examples of Platform Trials
2.1 A stylistic example
First, we present a stylistic example of a platform trial in Figure 1 to provide a concrete illustration of its key features. This trial comprises five sub-studies, each containing a randomized controlled comparison of several treatments. Initiating or stopping a sub-study is governed by a master protocol. Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study. Patients are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis. Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292).
Figure 1: A stylistic example of a platform trial. The x-axis represents calendar time discretized into windows with different potential enrollment in components of the platform trial. The figure illustrates both different treatment comparisons and different patient subgroups evolving over time.
In Figure 1, each box represents a sub-study along with its status, and each striped block within a box represents the enrollment period for a treatment evaluated in that sub-study. The horizontal axis denotes calendar time. Enrollment windows are defined by changes in the available treatments for enrollment. For instance, the initiation of enrollment for sub-study 3 marks the start of enrollment window 2, within which treatment 4 is compared to treatment 1 as the control among patients with disease subtype A. Likewise, the conclusion of the enrollment period for sub-study 2 defines the start of enrollment period 3. Subsequently, the start of treatment 3 as the new control in place of treatment 1 among subtype A defines the start of enrollment window 4.
The following is a narrative of Figure 1 according to the order of enrollment windows.
Enrollment Window 1. In this example, there are two initial sub-studies. Patients with any disease subtype is eligible to sub-study 1, while sub-study 2 is restricted to patients with subtype A. During enrollment window 1, patients with subtype A are randomly assigned to sub-study 1 or 2, and then randomized to treatment 1 or 2 if they are in sub-study 1 and to treatment 1 or 3 if they are in sub-study 2. Patients with disease subtypes other than A are enrolled in sub-study 1 and then randomized to treatment 1 or 2.
Enrollment Window 2. Sub-study 3 is added to the platform trial when a new intervention treatment 4 becomes available for disease subtype A. Following this addition, a patient with subtype A is randomly assigned to either sub-study 1, 2, or 3, and subsequently randomized to the treatments being assessed within the respective sub-study.
Enrollment Window 3. After sub-study 2 reaches full enrollment, it closes its enrollment. Later during this period, treatment 3 meets the pre-specified success criteria.
Enrollment Window 4. Since treatment 3 has been demonstrated to be superior to treatment 1 for subtype A, it replaces treatment 1 as the control for subtype A. Consequently, treatment 3 replaces treatment 1 in sub-study 3. Moreover, sub-study 1 excludes patients with subtype A, and sub-study 4 is introduced to continue evaluating treatment 2 in comparison to the new control treatment 3 for patients with subtype A.
Enrollment Window 5. Sub-study 5 is added upon the introduction of a new treatment 5 for patients positive for biomarker B. In sub-study 5, treatment 5 is evaluated against treatment 1 for patients without subtype A, or to treatment 3 for patients with subtype A.
Enrollment Window 6. Sub-study 4 has finished its planned enrollment and is no longer accepting new patients.
Enrollment Window 7. A new treatment, treatment 6, has been introduced in sub-study 5 for patients positive for biomarker B. During this enrollment window, treatment 2 in sub-study 4 fails to meet the pre-specified success criteria.
As the platform trial continues, additional sub-studies or treatments may be added as new interventions emerge for one or more targeted groups of patients, while existing sub-studies may be concluded due to either meeting success criteria or demonstrating futility.
As exemplified in Figure 1, platform trials can be very flexible: adding new sub-studies and/or treatments with different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial. Moreover, compared to a traditional clinical trial, it enables potential pooling of data across sub-studies. Specifically, in traditional clinical trials, patients self-select into studies and are randomized into treatment arms, with treatment comparisons conducted solely within each study. In contrast, platform trials enroll patients into a master protocol, randomizing them into eligible sub-studies and subsequently treatment arms, and consequently, patients can be pooled across sub-studies. For example, when comparing treatments 1 and 2, in a platform trial like Figure 1, patients from not only sub-study 1 but also other sub-studies receiving treatment 1 can be included in the analysis, whereas we analyze data within sub-study 1 in a traditional trial. See the simulation results and discussions in Section 7.1 for the efficiency gain.
2.2 A real example: the SIMPLIFY trial
Our second example is a real platform design aliased SIMPLIFY for people with cystic fibrosis (Mayer-Hamblett et al.,, 2023). It contains two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the Cystic Fibrosis Therapeutics Development Network. The primary objective for each study is to determine whether discontinuing either hypertonic saline (HS) or dornase alfa (DA), respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator, elexacaftor plus tezacaftor plus ivacaftor (ETI). The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1). Figure 2 depicts the SIMPLIFY study design.
Figure 2: SIMPLIFY study design. All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups: those receiving HS only, those receiving DA only, and those receiving both therapies. Individuals on a single therapy enter the corresponding study, while those on both are randomized to the HS or DA study, first with a 1:1 probability and then with a 3:1 probability during subsequent enrollment. All individuals in HS study (or DA study) are randomized 1:1 to continue or discontinue HS (or DA).
3 Setup and Estimand
3.1 Randomization procedures in platform trials
Consider a platform trial with $J$ treatments and $n$ patients. For each patient $i$, let $\bm{W}_{i}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A_{i}$ be the treatment assignment indicator that equals $j$ if the patient is assigned to arm $j$, $Y_{i}^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y_{i}=Y_{i}^{(A_{i})}$ be the observed outcome. Throughout, $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, are independent and identically distributed realizations of $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments. The following assumption describes the randomization procedure in platform trials.
Assumption 1 (Randomization).
There exists an observed baseline discrete variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
Assumption 1 highlights a fundamental difference between the randomization procedure used in platform trials versus traditional clinical trials. In traditional trials, the randomization probability is typically a fixed positive constant. However, in platform trials, this probability, $\pi_{j}(Z)$, varies depending on a baseline discrete variable $Z$ and may be zero for some treatments. Notably, $Z$ usually includes the enrollment window by calendar time and some key patient characteristics, as the enrollment window determines which treatments are available for enrollment, and patient characteristics determine the treatments for which a patient is eligible. Hence, $\pi_{j}(Z)$ may be zero if treatment $j$ has not entered the platform trial or has closed its enrollment, or if the patient does not meet the eligibility requirements for treatment $j$. Additionally, non-zero values of $\pi_{j}(Z)$ can differ over time and patient characteristics. Values of $\pi_{j}({Z})$, $j=1,...,J$, are known by design, but $\pi_{j}({Z})$ should be viewed as random due to the randomness of ${Z}$.
The example in Figure 1 illustrates these dynamics. Here, $Z=(Z_{\rm win},Z_{\rm sub},Z_{\rm B+})$ represent the enrollment window by calendar time, disease subtype (A or not A), and biomarker B (positive or negative), respectively, where $Z_{\rm B+}$ does not affect treatment assignment until enrollment window 5. For simplicity, consider randomization with equal probabilities. If patient $i$ with subtype A enters the trial during enrollment window 1 in which only sub-studies 1 and 2 are open, this patient is randomized to either sub-study with equal probability. If randomized to sub-study 1, the patient is further randomized to treatment 1 or 2 with equal probability. If randomized to sub-study 2, the patient is further randomized to treatment 1 or 3 with equal probability. Consequently, for this patient $i$, $\pi_{1}({Z}_{i})=0.5\times 0.5+0.5\times 0.5=0.5$, $\pi_{2}({Z}_{i})=0.5\times 0.5=0.25$, $\pi_{3}({Z}_{i})=0.5\times 0.5=0.25$, with all other $\pi_{j}({Z}_{i})$ values equal to 0. If another patient $i^{\prime}$ without subtype A enters the trial during the same enrollment window 1, then this patient is assigned to sub-study 1 with probability 1. Consequently, $\pi_{1}({Z}_{i^{\prime}})=\pi_{2}({Z}_{i^{\prime}})=0.5$ and all other $\pi_{j}({Z}_{i^{\prime}})$ values are 0. Similarly, if another patient $i{{}^{\prime\prime}}$ with subtype A enters the trial during enrollment window 2 in which sub-studies 1,2,3 are open, this patient has $\pi_{1}(Z_{i^{\prime\prime}})=0.5$, $\pi_{2}(Z_{i^{\prime\prime}})=\pi_{3}(Z_{i^{\prime\prime}})=\pi_{4}(Z_{i^{\prime\prime}})=1/6$, with all other $\pi_{j}({Z}_{i^{\prime\prime}})$ values equal to 0. These cases exemplify how the positivity and values of $\pi_{j}(Z)$ can vary with $Z$.
3.2 Treatment effects for concurrently eligible patients
We first define a clinically meaningful estimand (ICH E9 , 2019, R1) for comparing a given pair of treatments, $j$ and $k$, one of which is typically a control. As discussed in the introduction, many researchers and the FDA have recommended that the comparison be based on concurrently eligible patients, i.e., patients who meet the eligibility criteria for both treatments and are enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment $j$ or $k$. Formally, the concurrently eligible patients for treatments $j$ and $k$ are those with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
We provide two examples to illustrate this concept. The first example is based on Figure 1. When an investigator compares treatments 1 and 3 at the end of sub-study 2, the concurrently eligible patients are all subtype A patients from enrollment windows 1 and 2, which includes all patients in sub-study 2, as well as patients with subtype A in sub-study 1 and all patients in sub-study 3 from the first two windows. Patients without subtype A or enrolled in later windows are not included, as they are either ineligible or nonconcurrent. The second example is based on the SIMPLIFY design in Figure 2. The original analysis by Mayer-Hamblett et al., (2023) calculated the effects of discontinuing vs. continuing HS or DA within each study separately. However, this approach did not utilize all concurrently eligible patients with additional control data that could have been used from participants assigned to continue therapy across either study. Thus, when comparing discontinuing HS versus continuing therapy, for instance, the concurrently eligible patients are all individuals on HS regardless of their DA status, which includes all participants in the HS study, as well as those on both HS and DA in the DA study.
We propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ among concurrently eligible patients:
${\bm{\vartheta}}_{jk}=\left(\begin{array}[]{c}\theta_{jk}\\ \theta_{kj}\end{array}\right)=\left(\begin{array}[]{c}E(Y^{(j)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\\ E(Y^{(k)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\end{array}\right),$ (1)
where both subscripts $j$ and $k$ are used in $\bm{\vartheta}_{jk}$, $\theta_{jk}$, and $\theta_{kj}$ because both $\pi_{j}({Z})$ and $\pi_{k}({Z})$ are involved. In essence, a contrast of the estimand in (1) is an unconditional treatment effect for all concurrently eligible patients, averaging across all eligible patients and over time during which the enrollment window is open. Time can represent many background factors that may be changing although this is not explicitly recognized in estimand specification for traditional clinical trials.
Our defined population of concurrently eligible patients is the largest population that can preserve the integrity of randomization when inferring treatment effects. Individuals outside this group have zero probability of being assigned to treatments $j$ or $k$, or both, due to their ineligibility or the unavailability of treatments, thereby necessitating extrapolation. This is the reason why we do not consider the population means $E(Y^{(j)})$ and $E(Y^{(k)})$ that are defined for the entire platform trial population. Conversely, our estimand in (1) represents the treatment effect for a broader population compared to just those enrolled in a specific sub-study. Our estimand in (1) truly respects the eligibility criteria for the given pair of treatments for comparison, rather than on the subset assigned to a particular sub-study.
Throughout the paper, we use the notation convention that $a_{kj}$ is defined by swapping subscripts $j$ and $k$ in $a_{jk}$, e.g., $\theta_{jk}$ and $\theta_{kj}$ defined in (1). Based on the estimand ${\bm{\vartheta}}_{jk}$ in (1), the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$. Our approach to defining treatment effects is adaptable to situations where population characteristics other than means are considered, or when the treatment effect for a targeted subgroup is of primary interest.
4 Estimation Methods
4.1 Naive method
After we define the estimand ${\bm{\vartheta}}_{jk}$ in (1), the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible patients defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z}),\pi_{k}({Z})$ may still vary over time and patient characteristics, leading to confounding.
A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from patients with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$. While $\bar{Y}_{jk}$ is valid in traditional trials, it is typically biased under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome. Specifically, as shown in the Supplementary Material, as $n\to\infty$,
$\bar{Y}_{jk}\to\frac{E\{\pi_{j}({Z})E(Y^{(j)}\mid{Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}{E\{\pi_{j}({Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}\quad\mbox{in probability}.$ (2)
The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$. Take the trial described in Section 2.1 and Figure 1 as an example, where $\pi_{j}({Z})$ varies with ${Z}$, the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited patients being healthier than those recruited earlier. The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$. In traditional clinical trials, $\bar{Y}_{jk}$ is valid because $\pi_{j}({Z})$ does not vary with ${Z}$.
4.2 Inverse probability weighting
As noted in Section 4.1, the naive method is biased because it fails to account for confounding by $Z$. Under Assumption 1, since the true probability of receiving treatment $\pi_{j}(Z)$ is known, this confounding can be addressed using inverse probability weighting (IPW). The identification formula is (the proof is in the Supplementary Material):
$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},\vspace{-3mm}$ (3)
where $I(\cdot)$ denotes the indicator function. Formula (3) naturally suggests the following IPW estimator of $\theta_{jk}$, which is widely used in survey sampling (Horvitz and Thompson,, 1952) and causal inference (Rosenbaum and Rubin,, 1983),
$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},\vspace{-3mm}$
where ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ and $n_{jk}$ is the number of elements in ${{{\cal I}}}_{jk}$.
This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971; Robins et al.,, 2000) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
$\hat{\theta}_{jk}^{\rm\,sipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})}.$
Both the IPW and SIPW estimators are asymptotically normal (Theorem 1 of Section 5).
The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$. To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2023), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect. This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (ICH E9,, 1998; EMA,, 2015; FDA,, 2021).
Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may contain or overlap with ${Z}$. We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a potentially misspecified working model. For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y_{i}$ on $\bm{X}$ using data where $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$. For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model. After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for covariate $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aipw}=$ $\ \frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).\vspace{-3mm}$
Interestingly, in its last term, the AIPW estimator leverages covariate information from all concurrently eligible patients in ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$, including those not receiving treatments $j$ or $k$. For example, in Figure 1, when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2 using AIPW, the estimator also incorporates covariate data from subtype A patients assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2. These covariate data are not used in the IPW and SIPW estimators. Thus, AIPW offers a robust method to pool data from all concurrently eligible patients, thereby further extracting the potential of efficiency gain through platform trials.
Similarly, one can use the stabilized AIPW (SAIPW) estimator by normalizing the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$, resulting in the following SAIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,saipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
Both AIPW and SAIPW estimators are asymptotically normal (see Theorem 1 in Section 5). As discussed in Section 5, whether to stabilize the weights in the AIPW approach often results in asymptotically equivalent estimators.
4.3 Post-stratification
An alternative method to account for the varying $\pi_{j}({Z})$ is to stratify subjects based on the values that $\pi_{j}({Z})$ can take. This approach is known as post-stratification in sample surveys (Fuller,, 2009) or sub-classification in the causal inference literature (Rosenbaum and Rubin,, 1984). Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible patients ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$. The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
$\hat{\theta}_{jk}^{\rm\,ps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)Y_{i}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})\bar{Y}_{j}({\cal I}_{jk}^{(h)}),$
where $n({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$. This estimator does not need stabilizing, as the sum of weights is
$\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})=1.$
Compared with the SIPW, the PS requires an extra step of constructing strata for every pair $j$ and $k$.
The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time. For instance, in Figure 1, ${Z}$ includes both the disease subtype and the enrollment window; in the SIMPLIFY example in Figure 2, $Z$ is the indicator of HS and/or DA and there is only a single enrollment window. On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary. Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$. It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section 7.1.
The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$. To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages. This gives the following adjusted post-stratification (APS) estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{\cal I}}_{jk}^{(h)}}I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}+\frac{1}{n_{jk}}\sum_{i\in{{\cal I}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
Asymptotically, the PS and APS estimators are shown to be equivalent to certain types of SAIPW estimators, as detailed in Section 5.
4.4 Summary of estimators
[…]
5 Asymptotic Theory
5.1 Asymptotic normality
To develop the asymptotic theory for estimators of $\bm{\vartheta}_{jk}$ involving covariate adjustment, we need the following standard condition that the estimated working model $\hat{\mu}_{jk}$ converges to a well-defined limit $\mu_{jk}$, where the working model can be misspecified.
Assumption 2 (Stability).
For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$. If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in S1 of the Supplementary Material.
Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions. In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
Theorem 1.
Under Assumption 1, and Assumption 2 if adjusting for covariates, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$ whose explicit expression is given in each of the following specific case.
(a) For the
IPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ipw}$,
$\bm{\Sigma}_{jk}^{\rm\,ipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T},\vspace{-2mm}$
[…]
(b) For the
SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$,
$\bm{\Sigma}_{jk}^{\rm sipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)}-\theta_{jk})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)}-\theta_{kj})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}.\vspace{-2mm}$
(c) For the AIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aipw}$,
$\bm{\Sigma}_{jk}^{\rm aipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T},\vspace{-2mm}$
where $\bm{\delta}_{jk}=(\delta_{jk},\delta_{kj})^{T}$, $\delta_{jk}=E_{jk}\{Y^{(j)}\!-\mu_{jk}(\bm{X})\}$, $\mu_{jk}$ is the limit of $\hat{\mu}_{jk}$ given in Assumption 2, and $\bm{\Lambda}_{jk}$ is a symmetric matrix of order 2 with two diagonal elements $\lambda_{jk}=2\,{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and $\lambda_{kj}$, and the off-diagonal element $c_{jk}=c_{kj}=$ ${\rm Cov}_{jk}\{Y^{(j)},\mu_{kj}(\bm{X})\}+{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$.
(d) For the
SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$,
$\bm{\Sigma}_{jk}^{\rm saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})-\delta_{jk}\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})-\delta_{kj}\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}.$
(e) For the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$,
$\bm{\Sigma}^{\rm ps}_{jk}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}(Y^{(j)}\mid S)}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}(Y^{(k)}\mid S)}{\pi_{k}(S)}}\right\}\Big{]}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-2mm}$
where $S$ is the covariate for post-stratification, i.e., for patient $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplementary Material.
(f) For the
APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$,
$\bm{\Sigma}^{\rm aps}_{jk}$ $=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}\Big{]}$
$\quad+E_{jk}\{\bm{\Lambda}_{jk}(S)\}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-4mm}$
where $\bm{\Lambda}_{jk}(S)$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ replaced by $\lambda_{jk}(S)=2{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\mid S\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}$ and $c_{jk}(S)={\rm Cov}_{jk}\{Y^{(j)},$ $\mu_{kj}(\bm{X})\mid S\}$ $+\,{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\mid S\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}$, respectively.
All technical proofs are given in the Supplementary Material. The results in Theorem 1 for estimators with covariate adjustment hold regardless of whether the working model is correct or not. The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators in Section 6.
5.2 Efficiency comparison
As the six estimators proposed in Section 4 have different asymptotic covariance matrices, we compare the asymptotic relative efficiency between pairs of these estimators in the subsequent corollaries. The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
Corollary 1 (Comparison of IPW and SIPW).
Under Assumption 1, the difference in asymptotic covariance matrices of the IPW and SIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.$
Since the matrix $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ in Corolary 3 is positive semidefinite (shown in S6 of the Supplementary Material) but ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ may have arbitrary signs, the asymptotic relative efficiency of IPW versus SIPW is indefinite. When ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ are both non-negative, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ is positive semidefinite and, hence, SIPW always has equal or better asymptotic efficiency than IPW in estimating any differentiable function of $\bm{\vartheta}_{jk}$. On the other hand, an example of estimating $\theta_{jk}-\theta_{kj}$ is shown in S6 of the Supplement Material in which IPW is asymptotically more efficient than SIPW.
Nonetheless, SIPW is often more efficient in simulations and real data applications (see Section 7). SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change the SIPW estimator $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect the IPW analogue $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023). For these reasons, we do not consider IPW in the remaining efficiency comparisons.
Corollary 2 (Comparison of AIPW and SAIPW).
Under Assumptions 1-2, the difference in asymptotic covariance matrices of the AIPW and SAIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.\vspace{-2mm}$
Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition
$\delta_{jk}=\delta_{kj}=0,\vspace{-3mm}\vspace{-2mm}$ (4)
which is known as the prediction unbiasedness condition (Guo and Basse,, 2021) at the population level, where $\delta_{jk}=E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\}$ and $\delta_{kj}=E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\}$. Condition (4) holds when the working models used for covariate adjustment are correctly specified. It also holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified. Later discussions about adjustment (8) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet condition (4), it can be readily modified by (8) to satisfy this requirement. Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
The next two results show when covariate adjustment leads to an efficiency gain. Corollary 3 compares the two stabilized IPW estimators, with or without covariate adjustment.
Corollary 3 (Comparison of SIPW and SAIPW).
Under Assumptions 1-2 and
$E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}=\delta_{jk},\ \ E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}=\delta_{kj},$ (5)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{jk}(\bm{X})\mid S\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{kj}(\bm{X})\mid S\}=0,$ (6)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{kj}(\bm{X})\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{jk}(\bm{X})\}=0,$ (7)
where $S$ is the covariate for post-stratification as defined in Theorem 1(e), the SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ is asymptotically more efficient than the SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{\mu_{jk}(\bm{X})-\mu_{jk}\}^{2}}{\pi_{j}(S)},\,\frac{\{\mu_{kj}(\bm{X})-\mu_{kj}\}^{2}}{\pi_{k}(S)}}\right\}\Big{]}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$
is positive definite, where $\mu_{jk}=E_{jk}\{\mu_{jk}(\bm{X})\}$ and $\mu_{kj}=E_{jk}\{\mu_{kj}(\bm{X})\}$, unless either one of ${\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ is positive semidefinite.
When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect. Conditions (5)-(7) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the platform trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent. Actually, condition (5) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
$\tilde{\mu}_{jk}(\bm{X})=\hat{\mu}_{jk}(\bm{X})+\frac{1}{n_{jk}(S)}\sum_{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S}\{Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i})\},$ (8)
where $n_{jk}(S)$ is the number of elements in $\{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S\}$. The corresponding limit of $\tilde{\mu}_{jk}(\bm{X})$ in Assumption 2 is $\mu_{jk}(\bm{X})+E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}$ satisfying (5) with $\delta_{jk}=\delta_{kj}=0$. Conditions (6)-(7) hold when working models are correct. Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (6)-(7) embedded. For example, (5)-(7) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2023); see S10 in the Supplement Material for details.
In fact, (5)-(7) always hold for the SAIPW estimator with $\bm{X}=S$, i.e., the SAIPW estimator with $\hat{\mu}_{jk}(\bm{X})=\bar{Y}_{j}(S)$, which is the sample mean of $Y_{i}$’s for $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}=S$. This is because, with $S$ as the stratum indicator, the limit of $\bar{Y}_{j}(S)$ is $E_{jk}(Y^{(j)}\mid\bm{X}=S)$, which is always a correct working model, and consequently conditions (5)-(7) hold. Therefore, the SAIPW estimator with $\bm{X}=S$ is asymptotically more efficient than the SIPW estimator according to Corollary 3 and it is used in later corollaries to bridge the comparison between IPW-type and PS-type estimators.
Corollary 4 compares the two post-stratification estimators, with or without additional adjustment for covariates.
Corollary 4 (Comparison of PS and APS).
Under Assumptions 1-2 and (5)-(7), the APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,apw}$ is asymptotically more efficient than the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}\!=\!E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}\Big{]}$
is positive definite, unless either one of ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ conditioned on $S$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}$ is positive semidefinite.
Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in the trivial scenario where either no covariate adjustment beyond $S$ in one of the treatment arms or $\pi_{j}(S)+\pi_{k}(S)=1$ and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent conditioned on $S$.
To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting. For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$. This leads to the following result.
Corollary 5 (Comparison of PS and SAIPW with $\bm{X}=S$).
Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,ps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$ when $\bm{X}=S$.
Corollary 5 states that the PS estimator that stratifies on $S$ is asymptotically equivalent to the SAIPW estimator that adjusts for $S$ (as well as the corresponding AIPW estimator due to its equivalence with the SAIPW when (4) holds).
The next corollary compares APS and SAIPW with the same covariate adjustment. It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (5), the APS estimator is asymptotically equivalent to the SAIPW estimator.
Corollary 6 (Comparison of APS and SAIPW).
Under Assumptions 1-2 and condition (5), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,aps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$.
The last corollary completes the comparison around SAIPW with $\bm{X}=S$ and shows that under (5)-(7), adjusting for more covariates beyond just $S$ leads to additional asymptotic efficiency gain.
Corollary 7 (Comparison of SAIPW and SAIPW with $\bm{X}=S$).
Under Assumptions 1-2 and conditions (5)-(7), SAIPW with $S\subset\bm{X}$ is more efficient than SAIPW with $\bm{X}=S$.
All the efficiency comparisons are summarized in Figure 3.
Figure 3: Efficiency comparisons among all robust estimators.
6 Robust Variance Estimation
To robustly assess variability and conduct large sample inference, it is essential to obtain consistent estimators of the asymptotic covariance matrices in Theorem 1 for various estimators of $\bm{\vartheta}_{jk}$. These estimators should remain consistent whether the working models are correct or misspecified.
For IPW-type estimators, we propose the following estimators of covariance matrices:
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{\cal I}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)Y_{i}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)Y_{i}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}-\hat{\bm{\vartheta}}_{jk}\hat{\bm{\vartheta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,sipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)(Y_{i}-\hat{\theta}_{jk})^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)(Y_{i}-\hat{\theta}_{kj})^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk}-\hat{\bm{\delta}}_{jk}\hat{\bm{\delta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,saipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})-\hat{\delta}_{jk}\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})-\hat{\delta}_{kj}\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk},$
where $\hat{\bm{\vartheta}}_{jk}=(\hat{\theta}_{jk},\hat{\theta}_{kj})^{T}$, $\hat{\theta}_{jk}$ is any estimator of $\theta_{jk}$, $\hat{\bm{\delta}}_{jk}=(\hat{\delta}_{jk},\hat{\delta}_{kj})^{T}$, $\hat{\delta}_{jk}=n_{jk}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\!I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}/\pi_{j}({Z}_{i})$, $\hat{\bm{\Lambda}}_{jk}$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ respectively estimated by $\hat{\lambda}_{jk}=2\hat{q}_{jk}^{(j)}+\hat{\sigma}^{2}_{jk}$ and $\hat{c}_{jk}=\hat{q}_{kj}^{(j)}+\hat{q}_{jk}^{(k)}+\hat{q}_{jk}$, $\hat{q}_{jk}^{(j)}$ is the sample covariance of $(Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{jk}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$ and $A_{i}=j$, $\hat{q}_{kj}^{(j)}$ is $\hat{q}_{jk}^{(j)}$ with $\hat{\mu}_{jk}(\bm{X}_{i})$ switched to $\hat{\mu}_{kj}(\bm{X}_{i})$, $\hat{q}_{jk}^{(k)}$ is $\hat{q}_{jk}^{(j)}$ with $Y_{i}^{(j)}$ switched to $Y_{i}^{(k)}$, $\hat{\sigma}^{2}_{jk}$ and $\hat{\sigma}^{2}_{kj}$ are the diagonal entries and $\hat{q}_{jk}$ is the off-diagonal entry of the sample covariance matrix of $(\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{kj}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$.
For the PS and APS estimators, we propose the following estimators of $\bm{\Sigma}_{jk}^{\rm\,ps}$ and $\bm{\Sigma}_{jk}^{\rm\,aps}$,
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,{\rm diag}\!\left\{{\textstyle\frac{\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\sigma}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Gamma}}_{jk}$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,\Big{[}{\rm diag}\!\left\{{\textstyle\frac{\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\tau}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})\Big{]}+\hat{\bm{\Gamma}}_{jk},$
where $\hat{\pi}_{j}({\cal I}_{jk}^{(h)})=n_{j}({\cal I}_{jk}^{(h)})/n({\cal I}_{jk}^{(h)})$, $\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $Y_{i}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ estimated respectively by $\hat{\lambda}_{jk}({\cal I}_{jk}^{(h)})=2\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})+\hat{\sigma}_{jk}^{2}({\cal I}_{jk}^{(h)})$ and $\hat{c}_{jk}({\cal I}_{jk}^{(h)})=\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}^{(k)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}({\cal I}_{jk}^{(h)})$, $\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{\sigma}^{2}_{jk}({\cal I}_{jk}^{(h)})$, and $\hat{q}_{jk}({\cal I}_{jk}^{(h)})$ are respectively the same as $\hat{q}_{jk}^{(j)}$, $\hat{q}_{kj}^{(j)}$, $\hat{\sigma}^{2}_{jk}$, and $\hat{q}_{jk}$ previously defined in $\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}$ but with the index $i$ restricted to $i\in{\cal I}_{jk}^{(h)}$, $\hat{\bm{\Gamma}}_{jk}$ is the sample covariance matrix of $(\bar{Y}_{j}(S_{i}),\bar{Y}_{k}(S_{i}))$’s for $i\in{\cal I}_{jk}$, and $\bar{Y}_{j}(S_{i})$ is the sample mean of $Y_{i}$’s with $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}={\cal I}_{jk}^{(h)}$.
7 Empirical Results
7.1 Simulation study
We conduct a simulation study based on the first three enrollment windows outlined in Figure 1 to examine the finite-sample performance of all estimators presented in Section 4.
In the simulation, the observed baseline covariate vector is $\bm{W}=(X_{c},X_{b},Z_{\rm sub},Z_{\rm win})$, where $X_{c}$ is a continuous covariate uniformly distributed over the interval $(-3,3)$, $X_{b}$ is a binary covariate with $P(X_{b}=1)=0.5$, $Z_{\rm sub}$ is a binary indicator for disease subtype with $P(Z_{\rm sub}=1)=0.8$, and $Z_{\rm win}\in\{1,2,3\}$ denotes the enrollment window and satisfies
$P(Z_{\rm win}=t\mid X_{c},X_{b},Z_{\rm sub},U)=\frac{\exp(Q_{t})}{\exp(Q_{1})+\exp(Q_{2})+\exp(Q_{3})},\quad t=1,2,3,$
$Q_{1}=0.5+X_{c}+2X_{b}-Z_{\rm sub}+U$, $Q_{2}=1+2X_{c}+X_{b}-Z_{\rm sub}+U$, $Q_{3}=-0.5+X_{c}+X_{b}+Z_{\rm sub}+U$, and the unobserved baseline variable $U$ follows a standard normal distribution and is independent of $X_{c}$, $X_{b}$, and $Z_{\rm sub}$. The potential outcomes are generated as follows: $Y^{(1)}=1+X_{c}+X_{b}+Z_{\rm sub}+U+\epsilon_{1},$ $Y^{(2)}=1+X_{c}^{2}+X_{b}+Z_{\rm sub}+U+\epsilon_{2}$, $Y^{(3)}=3+X_{c}X_{b}+Z_{\rm sub}+U+\epsilon_{3}$, $Y^{(4)}=2+X_{c}Z_{\rm sub}-X_{b}+2U+\epsilon_{4}$, where $\epsilon_{j}$’s are mutually independent, each follows a standard normal distribution and is independent of $\bm{W}$ and $U$. Note that $Y^{(j)}$’s are correlated with $X_{c}$, $X_{b}$, $Z_{\rm sub}$, $U$, as well as indirectly with the enrollment time $Z_{\rm win}$.
Table 1: Assignment probabilities to each sub-study and treatment
${Z}$ Sub-study Treatment
$Z_{\rm win}$ $Z_{\rm sub}$ 1 2 3 $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$ $\pi_{4}({Z})$
1 1 0.4 0.6 0 0.50 0.20 0.30 0
1 0 1 0 0 0.50 0.50 0 0
2 1 0.3 0.3 0.4 0.50 0.15 0.15 0.20
2 0 1 0 0 0.50 0.50 0 0
3 1 0.4 0 0.6 0.50 0.20 0 0.30
3 0 1 0 0 0.50 0.50 0 0
Patients are first randomized into one of three sub-studies (labeled 1-3) depending on their enrollment windows and disease subtypes, and then to treatments within the assigned sub-study. Sub-study 1 has treatments 1 and 2, sub-study 2 has treatments 1 and 3, and sub-study 3 has treatments 1 and 4. Within each sub-study, treatments are randomized in a 1:1 ratio. The assignment probabilities are in Table 1. For the total sample size $n$ in the platform trial, we consider 500 and 1,000. When $n=500$, the expected numbers of patients per treatment are 123, 51, and 76 in sub-studies 1-3, respectively.
We consider the following estimators of the linear contrast $\theta_{j1}-\theta_{1j}$, for $j=2,3,4.$
  1. (i)
    Four IPW-type estimators. We consider the IPW, SIPW, and SAIPW estimators described in Section 4.2. For SAIPW, $\hat{\mu}_{jk}$ is obtained from fitting a linear regression of $Y$ on $(X_{c},X_{b},Z_{\rm sub})$, using data from treatment arm $j$ in ${\cal I}_{jk}$. Since condition (4) holds in this simulation, the AIPW estimator is numerically almost the same as the SAIPW estimator and thus is omitted. We add the SAIPW estimator with $\bm{X}=S$, denoted by SAIPW($S$), to check its equivalence with the PS estimator as discussed in Corollary 5. Note that the working models are correctly specified only for treatment arm 1, as the true models for the other arms are not linear in $(X_{c},X_{b},Z_{\rm sub})$.
  2. (ii)
    Four PS-type estimators. We consider the following strata for PS and APS:
    Estimating ${\small\begin{array}[]{ll}\mbox{Estimating $\theta_{21}-\theta_{12}$:}&{\cal I}_{12}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.5\,(Z_{{\rm sub},i}=0)\},\\ &{\cal I}_{12}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.2\,(Z_{{\rm win},i}=1\ \mbox{or}\ 3,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{12}^{(3)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{31}-\theta_{13}$:}&{\cal I}_{13}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.3\,(Z_{{\rm win},i}=1,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{13}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{41}-\theta_{14}$:}&{\cal I}_{14}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.2\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{14}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.3\,(Z_{{\rm win},i}=3,Z_{{\rm sub},i}=1)\}.\end{array}}\vspace{2mm}$
    The same working models $\hat{\mu}_{jk}$ for SAIPW are used for APS. For estimating $\theta_{21}-\theta_{12}$, we additionally include the PS and APS estimators that stratify on all six joint levels of ${Z}=(Z_{\rm win},Z_{\rm sub})$, denoted as PS($Z$) and APS($Z$). For estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$, we have PS($Z$) = PS and APS($Z$) = APS because using all six joint levels of ${Z}$ results in the same two strata as above (due to $\pi_{j}(Z)=0$ for some $j=3,4$).
  3. (iii)
    The naive estimator. For comparison we include the naive estimator of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, as described in Section 4.1.
  4. (iv)
    Estimators within each sub-study. For comparison, we also evaluate traditional ANOVA (sample means) and ANCOVA estimators (using linear working models to adjust for $X_{c},X_{b}$, and $Z_{\rm sub}$) based solely on data from each individual sub-study. Although these sub-study-specific estimators estimate the treatment effect within each sub-study, which differs from our target estimand defined in (1), we can still compare their variances.
The simulation results based on $5,000$ runs are reported in Table 2 for the previously described eleven estimators in (i)-(iv). The results contain the average bias, standard deviation (SD), the average of the standard error (SE), and the coverage probability (CP) of the $95\%$ confidence interval. The true values of estimands are approximated using an independent simulated dataset with size $10^{7}$. The following is a summary of the simulation results in Table 2.
Table 2: Simulation results based on $5,000$ simulation runs
$\theta_{21}-\theta_{12}=3$ $\theta_{31}-\theta_{13}=1.145$ $\theta_{41}-\theta_{14}=-0.886$
$n$ Method Bias SD SE CP Bias SD SE CP Bias SD SE CP
500 Naive -0.231 0.320 0.316 0.874 -0.185 0.342 0.340 0.916 -0.205 0.384 0.380 0.911
IPW -0.006 0.639 0.636 0.946 0.004 0.776 0.777 0.948 -0.007 0.500 0.497 0.948
SIPW -0.003 0.341 0.336 0.941 0.005 0.347 0.341 0.943 0.001 0.389 0.381 0.942
SAIPW -0.018 0.329 0.340 0.951 0.001 0.284 0.284 0.944 -0.001 0.297 0.300 0.949
$(S)$ 0.000 0.336 0.329 0.939 0.009 0.327 0.324 0.939 0.002 0.356 0.348 0.942
PS 0.000 0.336 0.335 0.945 0.009 0.327 0.330 0.949 0.002 0.356 0.356 0.946
APS -0.013 0.329 0.339 0.952 -0.001 0.286 0.289 0.947 -0.002 0.298 0.306 0.956
PS$(Z)$ Estimate or its SE cannot
APS$(Z)$ be computed in 408 runs
ANOVA* 0.001 0.354 0.350 0.946 0.006 0.421 0.417 0.945 0.004 0.424 0.425 0.946
ANCOVA* -0.008 0.329 0.321 0.942 -0.005 0.319 0.311 0.942 0.002 0.321 0.314 0.942
1000 Naive -0.230 0.226 0.224 0.819 -0.189 0.240 0.239 0.872 -0.206 0.269 0.268 0.876
IPW -0.001 0.453 0.451 0.947 0.012 0.550 0.550 0.951 0.003 0.355 0.352 0.943
SIPW 0.000 0.243 0.239 0.945 0.004 0.246 0.243 0.944 0.001 0.272 0.270 0.948
SAIPW -0.009 0.232 0.242 0.954 0.004 0.198 0.203 0.955 0.000 0.212 0.213 0.947
$(S)$ 0.001 0.238 0.234 0.945 0.004 0.233 0.230 0.942 0.003 0.252 0.247 0.943
PS 0.001 0.238 0.236 0.948 0.004 0.233 0.232 0.944 0.003 0.252 0.250 0.947
APS -0.006 0.232 0.239 0.952 0.003 0.198 0.203 0.955 0.000 0.213 0.215 0.952
PS$(Z)$ 0.001 0.238 0.236 0.948
APS$(Z)$0.005 0.233 0.229 0.942
ANOVA* 0.003 0.251 0.248 0.947 0.007 0.295 0.294 0.944 0.002 0.301 0.299 0.948
ANCOVA* -0.003 0.233 0.228 0.947 0.005 0.222 0.220 0.947 0.001 0.227 0.222 0.945
SAIPW($S$): SAIPW with $\bm{X}=S$.
PS($Z$) and APS($Z$): post-stratification using all joint levels of $Z=(Z_{\rm win},Z_{\rm sub})$ as strata (PS = PS($Z$) and APS = APS($Z$) for estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$).
*ANOVA and ANCOVA based on data within each sub-study, estimating the within sub-study treatment effects, which are $3.054$, $1.279$, and $-0.881$ for sub-studies 1-3, respectively.
  1. 1.
    The simulation results support our asymptotic theory in Section 5. Specifically, the naive estimator is biased, while all IPW-type and PS-type estimators have negligible biases; despite the misspecification of working models, covariate adjustment improves efficiency as reflected by reductions in SD; the PS and SAIPW($S$) estimators demonstrate nearly identical performance. Additionally, SEs are close to SDs and CPs are close to the nominal level 95%.
  2. 2.
    No results are reported in Table 1 for PS($Z$) and APS($Z$) at $n=500$ because either the estimate of $\theta_{jk}$ or its SE cannot be computed in 408 out of 5,000 simulation runs. This issue arises due to insufficient data in the stratum defined by $(Z_{\rm win},Z_{\rm sub})=(3,0)$, where the average number of patients is 10. A similar issue occurs in 3 out of 5,000 runs at $n=1,000$, so the results in Table 1 are based on 4,997 runs. This problem does not affect the PS and APS estimators: our proposed stratification method combines three strata $(Z_{\rm win},Z_{\rm sub})=(1,0),(2,0),(3,0)$ into ${\cal I}_{12}^{(1)}$ because they all have constant $\pi_{1}(Z)=0.5$ and $\pi_{2}(Z)=0.5$.
  3. 3.
    Between IPW and SIPW, it is clear that SIPW is much better in terms of SD under our simulation setting. Notably, IPW can have an even larger SD than ANOVA within substudies, which usually does not occur with SIPW.
  4. 4.
    The traditional ANOVA and ANCOVA estimators analyze the platform trials as if they were separate, stand-alone studies. Without covariate adjustment, the ANOVA estimator has a much larger SD than the SIPW estimator when estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$. A similar conclusion applies to the ANCOVA estimator when compared with the SAIPW or APS estimator. The efficiency difference is not as pronounced for estimating $\theta_{21}-\theta_{12}$, although the ANOVA and ANCOVA estimators still have larger SDs than the corresponding stabilized IPW-type and PS-type estimators. The reason for this is because the traditional ANOVA and ANCOVA do not utilize data across sub-studies, for example, when $n=500$, about 123, 51, and 76 patients are used under treatment arm 1 in sub-studies 1-3, respectively, whereas about 250, 121, and 150 patients in treatment arm 1 under platform trial are shared for our proposed estimators of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, respectively.
7.2 Real data analysis
As reviewed in Section 2.2 and shown in Figure 2, the SIMPLIFY design (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study. Let $Z$ be the indicator of whether a participant receives HS or DA therapy at baseline and the enrollment window, with the assignment probabilities given in Table 3. The enrollment window (EW) is treated as a categorical variable, with EW1 representing the enrollment period before the probability change, and EW2 standing for the enrollment period after the change, as presented in Figure 2. The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
In our analysis, the treatment effects of interest are $\theta_{21}-\theta_{12}$ and $\theta_{31}-\theta_{13}$. Specifically, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS among individuals taking HS, regardless of their DA status. This represents the treatment effect among participants with $\pi_{1}(Z)>0$ and $\pi_{2}(Z)>0$, pooled across studies (i.e., participants with $\text{HS}=1$ in Table 3). The interpretation applies similarly to $\theta_{31}-\theta_{13}$. Note that our estimand is slightly different from the estimand in Mayer-Hamblett et al., (2023), as the latter analyzed the HS and DA studies separately and included participants who re-enrolled for an additional study as noted in Section 2.2. We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
Table 3: Assignment probabilities in the SIMPLIFY trial
${Z}$ Study Treatment
HS DA EW HS DA $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$
1 1 EW1 0.50 0.50 0.50 0.25 0.25
1 1 EW2 0.75 0.25 0.50 0.375 0.125
1 0 any 1 0 0.50 0.50 0
0 1 any 0 1 0.50 0 0.50
We consider the IPW, SIPW, SAIPW, PS, and APS estimators, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section 7.1. For SAIPW and APS, in addition to $Z$, we adjust several other baseline covariates, including two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, delta F508 heterozygous, and other or unknown).
Estimates and 95% confidence intervals are presented in Figure 4, where all values are given in %. Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 4), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA. The conclusion is consistent with those in Mayer-Hamblett et al., (2023).
Figure 4: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial
8 Summary and Recommendations
Platform trials offer the potential to enhance statistical efficiency by facilitating the sharing of data between different treatment evaluations. However, realizing this potential requires careful consideration of several challenges, such as defining meaningful estimands that reflect eligibility and concurrency, and developing robust estimation techniques that accommodate the varying probabilities of treatment assignments over time and differing patient characteristics.
In this article, we explicitly present how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients. To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities. We also consider model-assisted approaches for covariate adjustment. We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID. Through theoretical analysis and simulation studies assessing the relative efficiency of these methods, we find that the stabilized IPW (SIPW) and post-stratification (PS) significantly enhance the efficiency of the IPW estimator. Additionally, model-assisted approaches for covariate adjustment (SAIPW or APS) generally improve efficiency while maintaining robustness against model misspecification. A detailed relative efficiency comparison is in Figure 3. Note that although SIPW is less efficient than PS asymptotically, it is still useful when there are sparse strata because SIPW may have better finite-sample performance and there may be situations where the PS estimator or its variance estimator cannot be computed.