### v4

# From Estimands to Robust Inference of Treatment Effects in Platform Trials

Abstract
A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments, where patients are often assigned to different subsets of treatment arms based on individual characteristics, enrollment timing, and treatment availability. While offering increased flexibility, this constrained and non-uniform treatment assignment poses inferential challenges, with two fundamental ones being the precise definition of treatment effects and robust, efficient inference on these effects. Such challenges arise primarily because some commonly used analysis approaches may target estimands defined on populations inadvertently depending on randomization ratios or trial operation format, thereby undermining interpretability. This article, for the first time, presents a formal framework for constructing a clinically meaningful estimand with precise specification of the population of interest. Specifically, the proposed entire concurrently eligible (ECE) population not only preserves the integrity of randomized comparisons but also remains invariant to both the randomization ratio and trial operation format. Then, we develop weighting and post-stratification methods to estimate treatment effects under the same minimal assumptions used in traditional randomized trials. We also consider model-assisted covariate adjustment to fully unlock the efficiency potential of platform trials while maintaining robustness against model misspecification. For all proposed estimators, we derive asymptotic distributions and propose robust variance estimators and compare them in theory and through simulations. The SIMPLIFY trial, a master protocol assessing continuation versus discontinuation of two common therapies in cystic fibrosis, is utilized to further highlight the practical significance of this research. All analyses are conducted using the R package RobinCID.
Keywords: Concurrently eligible individuals; Covariate adjustment; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
1 Introduction
1.1 Background
The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Woodcock and LaVange,, 2017). As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
Master protocol trials have emerged as innovative designs to accelerate the evaluation of potentially effective treatments (Woodcock and LaVange,, 2017). These trials evaluate one or more treatments across one or more diseases within an overarching study structure. Three major types of master protocol designs are umbrella trials, platform trials, and basket trials. Umbrella trials evaluate multiple therapies for a single disease (Rashdan and Gerber,, 2016; Renfro and Sargent,, 2017), basket trials evaluate a medical product for multiple diseases or disease subtypes (Redig and Jänne,, 2015; Hobbs et al.,, 2022), while platform trials offer the added flexibility of adding or removing treatment arms over time (Berry et al.,, 2015; Saville and Berry,, 2016; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023). These three types are not mutually exclusive, and a master protocol trial can incorporate features of more than one type. For example, an umbrella-platform hybrid for lung cancer might test different treatments across biomarker-defined subtypes (umbrella) while allowing treatments to enter or exit over time (platform). Similarly, a basket-platform hybrid might study a single targeted therapy across multiple cancers (basket), with new cancer types or biomarkers added adaptively over time (platform). Therefore, umbrella and basket trials can be viewed as special cases of platform trials, and, in this paper, we use the terms platform trial and master protocol interchangeably.
Master protocols originated in oncology and remain widely used in this field, where many new biomarker-driven therapies require efficient evaluation (Park et al.,, 2019, 2020; Ouma et al.,, 2022). By utilizing a common screening platform to efficiently screen for biomarkers and identify all treatments for which a patient is eligible, these trial designs significantly improve operational efficiency, accelerate patient accrual, and enable the study of rare tumor subtypes within one trial (Woodcock and LaVange,, 2017). In addition to streamlining operations, these designs also have the potential to improve statistical efficiency by facilitating data sharing across different treatment evaluations. Notable examples include the BATTLE umbrella trials (Kim et al.,, 2011; Papadimitrakopoulou et al.,, 2016), the Lung-MAP platform trial (Herbst et al.,, 2015), and the I-SPY 2 platform trial (Barker et al.,, 2009). The utility of master protocol trials is much broader. They are gaining prominence in other diseases areas, including infectious diseases (Normand,, 2021), neurodegenerative diseases such as amyotrophic lateral sclerosis (Quintana et al.,, 2023) and Alzheimer’s disease (Bateman et al.,, 2017), as well as inflammatory bowel disease (Honap et al.,, 2024). Section 2 presents two trials to provide concrete illustrations.
1.2 Two fundamental gaps
Compared to traditional parallel-group trials, master protocol trials often share a key feature: constrained and non-uniform treatment assignment. Constrained means some participants have zero probability of receiving certain treatments, while non-uniform refers to varying non-zero assignment probabilities based on participant characteristics, enrollment time, or treatment availability. This feature marks a key distinction between master protocol and traditional trial designs, and poses significant challenges for the interpretability of and statistical inference for the treatment effect, some of which are highlighted in a recent draft guidance from the U.S. Food and Drug Administration (FDA, 2023b, ). Although master protocols are increasingly conducted, two fundamental questions arising from their constrained and non-uniform treatment assignment remain unresolved.
1. How to define a clinically meaningful estimand?
An estimand is “a precise description of the treatment effect reflecting the clinical question posed by a given clinical trials objective” and is of ultimate importance (ICH E9 , 2019, R1). In master protocol trials, treatments often have different eligibility criteria and can be added or removed over time, leading to changes in the set of available treatment options throughout the trial. Consequently, the set of treatments to which an individual can be randomized and the corresponding assignment probabilities can vary over time and across individual characteristics. This leads to constrained and non-uniform treatment assignments. As a result, defining an estimand becomes more challenging because the treatment effect must be interpreted in the context of restricted eligibility, variable assignment probabilities, and evolving trial structure over time. In particular, among the five estimand attributes (ICH E9 , 2019 (R1)), the target population of interest requires special attention. Many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, 2023b have recommended using a population of individuals who are concurrently randomized and meet the eligibility criteria, as this preserves the integrity of randomized comparisons and avoids systematic differences between treatment groups due to measured and unmeasured factors. However, this general recommendation does not uniquely define the population, leading to ambiguity. For instance, in Cases I-II in Section 3.2, a seemingly reasonable population (and thus the estimand) may depend arbitrarily on randomization probabilities, even though it technically follows the general recommendation in FDA, 2023b .
To address this question, Section 3.2 introduces a population for comparing two treatments (typically with one as a control) that not only follows the general recommendations in FDA, 2023b but also respects the eligibility criteria while avoiding arbitrary dependencies on randomization probabilities. We refer to this as the entire concurrently eligible (ECE) population. This definition also applies to traditional parallel-group trials, to which we draw simple connections using Case I in Section 3.2 to explain why this population is sensible. The corresponding estimand for the ECE population is defined in Section 3.3. As in traditional trials, the target population in a master protocol trial may sometimes be a specific subgroup or a re-weighted population. In such cases, the ECE population can serve as the reference and may be restricted to the relevant subgroup or re-weighted to resemble the target population. Our proposal is the first in the literature to offer a systematic framework for defining clinically meaningful estimands with clear specification of the population of interest. Given the central importance of the estimand, our contribution enhances the fundamental understanding of master protocol trials and lays the groundwork for future statistical research—both frequentist and Bayesian—on topics such as interim analysis and the use of nonconcurrent controls.
2. How to provide robust estimation and inference regarding the defined treatment effect?
Although randomization is preserved, the commonly used estimator in traditional trials—comparing outcome means between treatment arms—can easily lead to biased estimates in master protocol trials, where treatment assignment is often constrained and non-uniform, potentially giving rise to issues such as Simpson’s paradox (Collignon et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021). Consequently, estimating treatment effects in master protocol trials is not straightforward.
A promising first step is to adapt inverse probability weighting or post-stratification to account for the constrained and non-uniform treatment assignment. Additionally, model-assisted covariate adjustment can offer another means to further unlock the efficiency potential of master protocol trials, while maintaining robustness against model misspecification. To guide practitioners in selecting among these methods, we derive the asymptotic variance and relative efficiency of these estimation methods. These results are novel and, to our knowledge, have not previously appeared in the literature.
1.3 Prior work
The increasing popularity of platform trials has led to a surge of research addressing their statistical challenges in both design and analysis. Many studies have focused on adaptive designs, which aim to efficiently screen candidate treatments and adaptively randomize individuals to the most promising treatments (Yuan et al.,, 2016; Kaizer et al.,, 2018; Ventz et al.,, 2018). Additionally, Bofill Roig et al., (2024) examined the optimal allocation ratios in these trials. A key concern in the analysis of platform trials is the potential presence of time trends, which can introduce bias when using nonconcurrent controls (Bofill Roig et al.,, 2023). To address this, various strategies have been proposed to account for time trends in the analysis (Elm et al.,, 2012; Saville et al.,, 2022; Bofill Roig et al.,, 2022; Jiang et al.,, 2023; Wang et al.,, 2023; Guo et al.,, 2024). However, all of these statistical methods have predominantly relied on correct modeling and are sensitive to model misspecification.
The platform trial literature on non-model-based approaches for concurrently eligible individuals remains limited. Marschner and Schou, (2022) proposed stratification by enrollment windows. Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that, conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows. More recently, Santacatterina et al., (2025) defined an estimand based on the concurrent period and established identification and estimation results using only concurrent controls. They also proposed strategies to incorporate nonconcurrent controls under a strong conditional exchangeability assumption between concurrent and nonconcurrent controls, given observed covariates. However, none of these studies consider the more general setting in which treatments have distinct eligibility criteria, nor do they rigorously investigate how to define a clinically meaningful estimand within a general framework that is independent of specific study designs or modeling assumptions.
1.4 Outline of the article
Section 2 presents the SIMPLIFY trial and a stylized trial used to emphasize key issues. Section 3 details the ECE population and corresponding estimand. In Section 4, we develop robust estimation methods, without or with covariate adjustment. In Section 5, we establish the asymptotic distributions and relative efficiency of the proposed estimators. We highlight the practical significance through the SIMPLIFY trial in Section 6, and we conclude with a summary and recommendations in Section 7.
2 Real and Stylistic Trials
2.1 The SIMPLIFY trial
We begin by presenting the real master protocol trial that motivates this work: the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023). Although it resembles an umbrella trial, where multiple treatments are evaluated concurrently and have different eligibility criteria, it also presents a key feature of platform trials: varying randomization ratio over time. Despite its seemingly simple design diagram, the trial involves both constrained and non-uniform treatment assignments. Its clarity makes it ideal for highlighting the two fundamental gaps, as it facilitates understanding of the underlying challenges.
Cystic fibrosis (CF) is caused by mutations in the CF transmembrane conductance regulator (CFTR) gene. CFTR modulator therapies are a class of small-molecule drugs designed to address the underlying cause of CF, leading to transformative improvements in patient outcomes. A landmark example is the triple-combination therapy elexacaftor/tezacaftor/ivacaftor (ETI), approved for individuals with at least one F508del allele, a mutation present in over 85% of people with CF. Given the potential overlap in therapeutic effects between ETI and pre-existing treatments such as hypertonic saline (HS) and dornase alfa (DA) and the need to reduce treatment burden, it is important to evaluate the impact of discontinuing these pre-existing therapies in individuals taking ETI. The SIMPLIFY trial was conducted within this context. It includes two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the CF Therapeutics Development Network. The primary objective for each study is to determine whether discontinuing either HS or DA, respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator ETI. The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1).
Figure 1: SIMPLIFY design schema. Individuals on a single therapy enter into the corresponding study with probability 1, while those on both are randomized to the HS or DA study, with ratio of 1:1 during enrollment window (EW) 1, then 3:1 during EW 2. Within HS (or DA) study, individuals are randomized 1:1 to continue or discontinue HS (or DA).
Table 1: Assignment probabilities in the SIMPLIFY trial.
Baseline variables EW Studies Treatments
HS DA EW HS study DA study 1 2 3
1 1 EW1 0.50 0.50 0.50 0.25 0.25
1 1 EW2 0.75 0.25 0.50 0.375 0.125
1 0 any 1 0 0.50 0.50 0
0 1 any 0 1 0.50 0 0.50
Figure 1 depicts the SIMPLIFY study design. All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups based on medication use when entering the trial: those on HS only, DA only, or both therapies. Individuals on a single therapy (HS only or DA only) are assigned to the corresponding study with probability 1. Those on both therapies are randomized to the HS or DA study, with 1:1 ratio during initial enrollment window 1 (before January 14, 2022), and 3:1 ratio during enrollment window 2 (after January 14, 2022) to assign more participants to the HS study, as the DA study was enrolling more quickly. This swift shift in the randomization ratio exemplifies how the flexibility of master protocol trials can accelerate clinical trials. Within each study (HS or DA), participants were then randomized 1:1 to either continue or discontinue their respective therapy. Table 1 summarizes assignment probabilities of the SIMPLIFY trial. It illustrates both constrained and non-uniform treatment assignment: participants receiving only HS or DA have zero probability of being assigned to certain treatment arms, and the non-zero probabilities of receiving any given pair of treatments may vary based on baseline characteristics and enrollment windows. Detailed Analysis is carried out in Section 6.
2.2 A stylistic trial
To motivate a general statistical framework for master protocol trials, we introduce a stylistic trial in two operation formats, both can be unified under our proposed framework. This unification highlights an important point: while trial design may influence clinical operations, it should not alter the clinical questions being addressed. Figure 2(a) presents the stylistic trial in a sub-study format, where treatments are evaluated in separate sub-studies. Figure 2(b) shows a multi-arm format, where each treatment is studied in a separate arm without dividing the trial into sub-studies.
The master protocol trial (in sub-study format) in Figure 2(a) comprises four sub-studies, each containing a randomized controlled comparison of two treatments. Initiating or stopping a sub-study is governed by a master protocol. Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study. Individuals are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis. Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292). A narrative of Figure 2(a) is included in Section S1 of the Supplement.
Figure 2: A stylistic master protocol trial in two operating formats: (a) sub-study format, (b) multi-arm format. Each box in (a) represents a sub-study, along with its eligibility criteria and status. Each striped block within a box in (a) or standalone in (b) represents the enrollment period for a treatment evaluated. The horizontal axis shows calendar time, divided into discrete windows based on changes in available treatments for enrollment.
In Figure 2(b), we reframe the same master protocol trial into a multi-arm format. In enrollment window 1, there are three initial treatment arms: investigational treatments 2 and 3 and treatment 1 serving as the common control. In this window, individuals with subtype A can be randomized to any of the three arms, while those with subtypes other than A can only be randomized to treatments 1 or 2. Descriptions for subsequent enrollment windows are omitted due to their similarity with Figure 2(a). Examples of real trials that follow this structure include the STAMPEDE trial for prostate cancer (Sydes et al.,, 2012) and the RECOVERY trial for COVID-19 therapies (Normand,, 2021).
As exemplified in Figure 2, master protocol trials can be very flexible: adding new sub-studies and/or treatments with new interventions and different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial.
3 Estimand
3.1 Randomization procedure
Consider a platform trial with $J$ treatments. For each individual, let $\bm{W}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A$ be the treatment assignment indicator that equals $j$ if the individual is assigned to arm $j$, $Y^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y=Y^{(A)}$ be the observed outcome. We assume throughout that $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, is an independent and identically distributed sample of size $n$ from $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments. The following is assumed for randomization in platform trials.
Assumption 1 (Constrained and non-uniform randomization).
There exists an observed baseline variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
Assumption 1 highlights a fundamental difference between platform and traditional trials. Traditional trials typically use a fixed randomization probability, whereas in platform trials, the randomization probabilities, $\pi_{j}(Z)$, vary with a baseline variable $Z$, which usually is discrete and may include enrollment window (defined by calendar time), site, and individual characteristics. $\pi_{j}(Z)$ may even be zero for some treatments. For example, if treatment $j$ is not yet available, has closed enrollment, is not offered at a given site, or the individual is ineligible. Even when non-zero, $\pi_{j}(Z)$ can vary across individuals and over time. Although $\pi_{j}({Z})$ is known by design, it is random due to the randomness of ${Z}$.
[…]
3.2 Population
Building on the discussion in Section 1.2, we now formally define the population of interest for comparing two treatments $j$ and $k$ (one of which is typically a control). In this paper, population specifically refers to the population of individuals used to define the estimand.
Definition (ECE population).
For the given treatments $j$ and $k$, the entire concurrently eligible (ECE) population is a population of all individuals who meet the eligibility criteria for both treatments and could potentially be enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment. Formally, the ECE population is a population of all individuals with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
The ECE population is a population of all individuals who could potentially be assigned to treatments $j$ and $k$, rather than the population represented by those actually assigned to these treatments (i.e., the population represented by $A\in\{j,k\}$). In traditional trials, this distinction is often overlooked because the potential and actual populations coincide. However, in platform trials, this issue has been brought to the forefront.
We present two basic cases to illustrate why using the population represented by those actually assigned to the two treatments or sub-studies may lead to unexpected problems. In Case I, all participants are eligible for all treatments, and the focus is on how changing treatment allocation probabilities over time (non-uniform treatment assignment) can cause issues in defining the population. In Case II, all treatments enroll concurrently, and the focus is on how different eligibility criteria for different treatments (constrained treatment assignment) complicates population definition. Finally, we revisit the SIMPLIFY trial, which features both constrained and non-uniform treatment assignment, reflected in time-varying treatment allocation probabilities and varying eligibility criteria across treatments.
Figure 3: Comparison of traditional parallel-group and platform trial designs in a hypothetical mild to moderate disease population (Case I). In this case, the ECE population for comparing treatments 2 and 1 coincides with the entire trial population across all designs. In contrast, in platform trials, the population represented by those assigned to specific treatments or sub-studies may vary arbitrarily depending on the trial design or randomization probabilities.
Case I in Figure 3 shows various trial designs evaluating treatments 1 and 2, with all participants eligible for all treatments. The entire trial population is 50% mild and 50% moderate, but patient characteristics shift over time: enrollment window 1 representing one third of the population consists of 30% mild and 70% moderate, while enrollment window 2 comprising the remaining two third includes 60% mild and 40% moderate. We focus on the population for comparing treatments 2 and 1, as these two treatments are open for enrollment during the same period across all designs.
If a traditional parallel-group trial design (Design 1) is used, all participants have a positive probability of receiving either treatment, making the ECE population identical to the entire trial population (50% mild and 50% moderate). This is typically the default population for comparing treatments 1 and 2. The same applies to all platform trial designs in Figure 3, where the ECE population remains identical to the entire trial population regardless of the specific platform trial type (Design 2 vs. 3) or randomization probabilities (Design 2 vs. 2’). This consistency provides a principled justification for using the ECE population, as the target population for treatment comparisons should not depend on the specific trial operation formate or randomization probabilities, in the same way that choices of randomization ratio do not affect the study population in traditional trials.
However, in platform trials, it is common to conduct analyses that naively subset participants assigned to treatment 1 or 2 (Design 2 and 2’), or to sub-study 1 (Design 3). Even without explicit subsetting, methods such as the t-test are often applied in Designs 2 and 2’, where the randomization ratio between treatments 1 and 2 is constant. These analyses target the population of individuals assigned to specific treatments or sub-studies. Although such analysis populations still consist of concurrently eligible individuals, as shown in Figure 3, they differ from the ECE population and can be unexpectedly influenced by design formats and randomization ratios. See Section S1 of the Supplement for more discussions.
Lastly, we remark that the practice of focusing on the analysis populations assigned to specific treatments or sub-studies likely originates from traditional parallel-group trials. In traditional trials, these issues do not arise because the constant randomization ratio ensures that the analysis population in any arm coincides with the ECE population.
Figure 4: A hypothetical umbrella trial (Case II) with a population with different disease severity (mild, moderate, and severe). In this case, the ECE population for comparing treatments 2 and 1 is the mild-to-moderate disease population (with 50% mild and 50% moderate). In contrast, the analysis population defined by participants assigned to sub-study 1 may vary arbitrarily depending on how individuals with moderate disease are randomized between the two sub-studies.
Case II considers an umbrella trial where all treatments enroll concurrently. There are two sub-studies: one evaluates treatments 1 and 2 in a population with mild to moderate disease, and the other evaluates treatments 3 and 4 in a population with moderate to severe disease. The moderate disease population is randomized between the two sub-studies with a 1:1 ratio. These two populations are commonly of interest in the context of ulcerative colitis (FDA,, 2022). Suppose the overall disease population consists of 40% mild, 40% moderate, and 20% severe disease.
When studying treatments for mild-to-moderate disease, the target population is 50% mild and 50% moderate. As shown in Figure 4, the ECE population for comparing treatments 1 and 2 includes all individuals with mild or moderate disease, since all can potentially be assigned to these two treatments, which exactly represents the population of interest – 50% mild and 50% moderate. In contrast, if only considering those participants assigned to treatments 1 or 2, or those within sub-study 1, the analysis population may arbitrarily depend on the randomization ratio between sub-studies. For instance, if the allocation ratio to sub-study 1 for the moderate disease population decreases from 1/2 to 1/4, the resulting population restricted to sub-study 1 shifts to 80% mild and 20% moderate.
The SIMPLIFY trial has two sub-studies: one compares treatments 1 and 2 in a population on HS only or both HS and DA; the other compares treatments 1 and 3 in a population on DA only or both HS and DA. Participants on both HS and DA are first randomized between the two sub-studies using a 1:1 ratio, which was later adjusted to 3:1. As shown in Figure 1, the ECE population for comparing treatments 1 and 2 (i.e., assessing the effect of discontinuing HS versus continuing) includes all individuals on HS only or on both HS and DA, irrespective of the sub-study randomization ratio. This appropriately reflects the eligibility criteria for the treatments being compared. In contrast, a naive analysis restricted to participants assigned to the HS study targets an analysis population that differentially weights the HS-only and HS+DA groups, where the weights are determined by the randomization ratio among individuals on both HS and DA, as well as the timing of any changes to that ratio; further discussions are in section 6. Alternatively, an analysis that includes all participants assigned to treatments 1 or 2 may include DA-only individuals, making the resulting analysis population less relevant for evaluating the effect of discontinuing HS.
In summary, the ECE population preserves the integrity of randomized comparison, is invariant to the randomization ratio or trial type, and aligns with the intention-to-treat principle (ICH E9,, 1998). It is an appropriate population for both traditional and platform trials, and is especially useful for providing a clear and logical framework for defining populations when evaluating treatment effects in platform trials. In contrast, any population defined by actual treatment assignments or sub-study allocations may not be reasonable, as they can be quite arbitrary and influenced by trial design or specific randomization probabilities. Additionally, individuals outside the ECE population have zero probability of being assigned to either treatment $j$ or $k$, due to ineligibility or unavailability of treatments, thereby necessitating extrapolation.
In some situations, other populations, such as subgroups defined by specific characteristics, enrollment windows, or re-weighted populations from different enrollment windows, may be of interest. For example, in COVID-19 trials with rapidly evolving variants, the population enrolled during later windows, when newer variants are prevalent, may hold the greatest clinical relevance. The ECE population weights each enrollment window according to its natural size, but alternative weights that reflect clinical priorities can also be considered. The concept of ECE population can also be extended to comparisons involving more than two arms. For instance, the ECE population for comparing three treatments $j$, $k$, and $l$ is a population of all individuals for whom $\pi_{j}(Z)>0$, $\pi_{k}(Z)>0$, and $\pi_{l}(Z)>0$. In any case, we recommend starting with the ECE population as the baseline, then considering its subgroups or applying weighting to reflect the target population of interest. This approach offers clarity and specificity in defining the population of interest, ensuring that the population reflects the clinical question rather than having arbitrary dependence on design choices like randomization ratio or trial types. Moreover, the ECE population is crucial in other areas of statistical research, such as when using nonconcurrent controls to improve trial efficiency, as it provides a clear reference point for evaluating both efficiency gains and potential bias. These topics will be explored in future research.
3.3 Treatment effect for the ECE population
With the population now clearly defined, we propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ within the ECE population:
${\bm{\vartheta}}_{jk}=\left(\begin{array}[]{c}\theta_{jk}\\ \theta_{kj}\end{array}\right)=\left(\begin{array}[]{c}E(Y^{(j)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\\ E(Y^{(k)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\end{array}\right),$ (1)
where $\theta_{jk}$ and $\theta_{kj}$ depend on $j$ and $k$ because the ECE population depends on them. A contrast of the estimand in (1) is an unconditional treatment effect for the ECE population, averaging across all eligible participants and over time during which the enrollment window is open. For example, the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$. In the context of the SIMPLIFY trial, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS versus continuing in a population of all individuals taking only HS or both HS and DA. A similar interpretation applies to $\theta_{31}-\theta_{13}$. This estimand framework is readily extendable to settings with multiple treatment arms.
Due to differences in the populations, $\theta_{jk}$ and $\theta_{kj}$ in (1) differ from $E(Y^{(j)})$ and $E(Y^{(k)})$, respectively, which are the means for the entire master protocol trial population and may include individuals who are non-concurrent or ineligible for treatments $j$ and $k$. Inference on $E(Y^{(j)})$ and $E(Y^{(k)})$ requires extrapolation. When these quantities are clinically of interest, one can apply methods to generalize the estimand in (1) to the entire trial population under additional assumptions (Degtiar and Rose,, 2023). Additionally, the estimand in (1) differs from the treatment effect for those assigned to specific arms (e.g., $E\{Y^{(j)}-Y^{(k)}\mid A\in(j,k)\}$), which, as shown by the formula, can depend on the distribution of $A$ and have problems such as arbitrary dependence on the randomization ratio. Similarly, the estimand in (1) also differs from the treatment effect for those assigned to specific sub-studies.
4 Estimation Methods
4.1 Analysis set
The analysis set, distinct from the population used to define the estimand, refers to the sample used in statistical analysis. While the ECE population is uniquely specified, the analysis set is not unique and can include multiple options. To maximize efficiency, the analysis set can include all concurrently eligible individuals in the index set ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$. However, practical considerations may lead to using a subset of ${{{\cal I}}}_{jk}$. For example, when applying a partial blinding strategy (FDA, 2023b, ) or if other sub-studies remain blinded at the time of analysis, it may be preferable to restrict the analysis set to individuals within a specific sub-study. In these cases, the analysis set is a biased sample of the ECE population and requires appropriate corrections. A general discussion of analysis sets and the conditions they must satisfy is provided in Section S1 of the Supplement. To focus our discussions, in the following, we consider either the analysis set defined by ${{{\cal I}}}_{jk}$ (with covariate adjustment) or the subset ${{{\cal I}}}_{jk}\cap\{i:A_{i}=j\text{ or }A_{i}=k\}$ (without covariate adjustment), where $\cap$ denotes the intersection of two sets.
4.2 Naive method
After we define the estimand, the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible individuals defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z})$ and $\pi_{k}({Z})$ may still vary over time and individual characteristics, leading to confounding.
A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from individuals with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$. However, $\bar{Y}_{jk}$ is typically biased for $\theta_{jk}$ under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome. Specifically, as shown in the Supplement, as $n\to\infty$,
$\bar{Y}_{jk}\to\frac{E\{\pi_{j}({Z})E(Y^{(j)}\mid{Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}{E\{\pi_{j}({Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}\quad\mbox{in probability}.$ (2)
The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$. Take the trial described in Figure 2 as an example, where $\pi_{j}({Z})$ varies with ${Z}$ (which does not occur in traditional clinical trials), the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited individuals being healthier than those recruited earlier. The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
4.3 Inverse probability weighting
As noted in Section 4.2, the naive approach is biased because it does not account for confounding by $Z$. Under Assumption 1, since the true probability of receiving treatment, $\pi_{j}(Z)$, is known, this confounding can be addressed through inverse probability weighting (IPW). Notably, weighting by $\pi_{j}(Z)$ also helps correct for sample bias, as IPW uses only those assigned to treatment $j$, which might otherwise be a biased sample from the ECE population in platform trials (see the discussion at the end of Section 3.2).
The IPW identification formula is (the proof is in the Supplement):
$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},$ (3)
where $I(\cdot)$ denotes the indicator function. Formula (3) naturally suggests the following IPW estimator of $\theta_{jk}$, which is widely used in survey sampling (Horvitz and Thompson,, 1952) and causal inference (Rosenbaum and Rubin,, 1983),
$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},$
where ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ and $n_{jk}$ is the number of elements in ${{{\cal I}}}_{jk}$.
This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
$\hat{\theta}_{jk}^{\rm\,sipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})}.$
Both the IPW and SIPW estimators are asymptotically normal (Theorem 1 of Section 5).
The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$. To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2025), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect. This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (FDA, 2023a, ).
Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may overlap with ${Z}$. We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a working model that may be misspecified. For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y^{(j)}$ on $\bm{X}$ using data with $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$. For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model. After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aipw}=$ $\ \frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).\vspace{-3mm}$
Interestingly, in its last term, the AIPW estimator leverages covariate information from all individuals within ${{{\cal I}}}_{jk}$, including those not receiving treatments $j$ or $k$. For example, in Figure 2(a), when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2, the AIPW also incorporates covariate data from subtype A individuals assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2. These covariate data are not used in the IPW or SIPW estimators. Thus, AIPW offers a robust method to pool data from all individuals within ${{{\cal I}}}_{jk}$, thereby further extracting the potential of efficiency gain through platform trials.
Similarly, one can use the stabilized AIPW (SAIPW) estimator by normalizing the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$, resulting in the following SAIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,saipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
Both AIPW and SAIPW estimators are asymptotically normal (Theorem 1 in Section 5). Note that when $Z_{i}$ is discrete, $\pi_{j}(Z_{i})$ can be estimated using sample proportions and used in place of the true value $\pi_{j}(Z)$ in any of the above weighting estimators.
4.4 Post-stratification
When $\pi_{j}({Z})$ takes discrete values, an alternative method to account for the varying $\pi_{j}({Z})$ is to stratify individuals based on the values that $\pi_{j}({Z})$ can take. This approach is known as post-stratification in sample surveys (Fuller,, 2009). Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible individuals in ${{{\cal I}}}_{jk}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$. The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
$\hat{\theta}_{jk}^{\rm\,ps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)Y_{i}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})\bar{Y}_{j}({\cal I}_{jk}^{(h)}),$
where $n({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$. This estimator does not need stabilizing, as the sum of weights in $\hat{\theta}_{jk}^{\rm\,ps}$ is equal to 1.
The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time. On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary. Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$. It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section S2 of the Supplement.
The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$. To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages. This gives the adjusted post-stratification (APS) estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{\cal I}}_{jk}^{(h)}}I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}+\frac{1}{n_{jk}}\sum_{i\in{{\cal I}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
5 Asymptotic Theory
5.1 Asymptotic normality
To develop the asymptotic theory for estimators of $\bm{\vartheta}_{jk}$ involving covariate adjustment, we need the following standard condition that the estimated working model $\hat{\mu}_{jk}$ converges to a well-defined limit $\mu_{jk}$, where the working model can be misspecified.
Assumption 2 (Stability).
For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$. If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in Section S1 of the Supplement.
Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions. In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$. Similarly, $E_{jk}(\cdot\mid S)=E_{kj}(\cdot\mid S)$, ${\rm Var}_{jk}(\cdot\mid S)={\rm Var}_{kj}(\cdot\mid S)$, and ${\rm Cov}_{jk}(\cdot\mid S)={\rm Cov}_{kj}(\cdot\mid S)$ denote these quantities further conditioned on $S$, where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$.
Theorem 1.
Under Assumption 1, and Assumption 2 if covariates are adjusted, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$, where $\hat{\bm{\vartheta}}_{jk}^{\star}=(\hat{\theta}_{jk}^{\star},\hat{\theta}_{kj}^{\star})^{T}$, $\bm{a}^{T}$ is the vector transpose of $\bm{a}$, and the explicit expression of $\bm{\Sigma}^{\star}_{jk}$ is given in each of the following specific case.
(a) For the
IPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ipw}$,
$\bm{\Sigma}_{jk}^{\rm\,ipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T},\vspace{-2mm}$
[…]
(b) For the
SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$,
$\bm{\Sigma}_{jk}^{\rm sipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)}-\theta_{jk})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)}-\theta_{kj})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}.\vspace{-2mm}$
(c) For the AIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aipw}$,
$\bm{\Sigma}_{jk}^{\rm aipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T},\vspace{-2mm}$
where $\bm{\delta}_{jk}=(\delta_{jk},\delta_{kj})^{T}$, $\delta_{jk}=E_{jk}\{Y^{(j)}\!-\mu_{jk}(\bm{X})\}$, $\mu_{jk}$ is the limit of $\hat{\mu}_{jk}$ given in Assumption 2, and $\bm{\Lambda}_{jk}$ is a symmetric matrix of order 2 with two diagonal elements $\lambda_{jk}=2\,{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and $\lambda_{kj}$, and the off-diagonal element $c_{jk}=c_{kj}=$ ${\rm Cov}_{jk}\{Y^{(j)},\mu_{kj}(\bm{X})\}+{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$.
(d) For the
SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$,
$\bm{\Sigma}_{jk}^{\rm saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})-\delta_{jk}\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})-\delta_{kj}\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}.$
(e) For the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$,
$\bm{\Sigma}^{\rm ps}_{jk}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}(Y^{(j)}\mid S)}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}(Y^{(k)}\mid S)}{\pi_{k}(S)}}\right\}\Big{]}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-2mm}$
where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplement.
(f) For the
APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$,
$\bm{\Sigma}^{\rm aps}_{jk}$ $=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}\Big{]}$
$\quad+E_{jk}\{\bm{\Lambda}_{jk}(S)\}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-4mm}$
where $\bm{\Lambda}_{jk}(S)$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ replaced by $\lambda_{jk}(S)=2{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\mid S\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}$ and $c_{jk}(S)={\rm Cov}_{jk}\{Y^{(j)},$ $\mu_{kj}(\bm{X})\mid S\}$ $+\,{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\mid S\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}$, respectively.
All technical proofs are given in the Supplement. The results in Theorem 1 and the efficiency comparison in Section 5.2 are new to the literature and differ from existing results for IPW or AIPW estimators using alternative weights, such as estimated propensity scores (Robins et al.,, 1994; Lunceford and Davidian,, 2004). We also show that the AIPW and SAIPW estimators with $\mu_{jk}=E_{jk}(Y^{(j)}\mid\bm{X})$ and $\mu_{kj}=E_{jk}(Y^{(k)}\mid\bm{X})$ are asymptotically equivalent and both attain semiparametric efficiency. The proof is provided in Section S3 of the Supplement. Furthermore, our results include PS and APS estimators and hold regardless of whether the working model is correctly specified. The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators presented in Section S1 of the Supplement.
5.2 Efficiency comparison
We compare the asymptotic relative efficiency between pairs of six estimators in the subsequent corollaries. The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
Corollary 1 (Comparison of IPW and SIPW).
Under Assumption 1, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
Although $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ is positive semidefinite, the asymptotic relative efficiency of IPW versus SIPW is indefinite because the covariances may have arbitrary signs; see Section S3 of the Supplement for more discussions. Nonetheless, SIPW is often more efficient in simulations (see Section S2 of the Supplement). SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023). For these reasons, we do not consider IPW in the remaining efficiency comparisons.
Corollary 2 (Comparison of AIPW and SAIPW).
Under Assumptions 1-2, $\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},}$ ${\textstyle{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition $\delta_{jk}=\delta_{kj}=0$, which holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified. Later discussions about adjustment (7) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet this condition, it can be readily modified by (7) to satisfy this requirement. Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
The next two results show when covariate adjustment leads to an efficiency gain.
Corollary 3 (Comparison of SIPW and SAIPW).
Under Assumptions 1-2 and
$E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}=\delta_{jk},\ \ E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}=\delta_{kj},$ (4)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{jk}(\bm{X})\mid S\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{kj}(\bm{X})\mid S\}=0,$ (5)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{kj}(\bm{X})\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{jk}(\bm{X})\}=0,\vspace{-2mm}$ (6)
where $S$ is the covariate for post-stratification as defined in Theorem 1(e), the SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ is asymptotically more efficient than the SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{\mu_{jk}(\bm{X})-\mu_{jk}\}^{2}}{\pi_{j}(S)},\,\frac{\{\mu_{kj}(\bm{X})-\mu_{kj}\}^{2}}{\pi_{k}(S)}}\right\}\Big{]}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$
is positive definite, where $\mu_{jk}=E_{jk}\{\mu_{jk}(\bm{X})\}$ and $\mu_{kj}=E_{jk}\{\mu_{kj}(\bm{X})\}$, unless either one of ${\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ is positive semidefinite.
When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect. Conditions (4)-(6) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the master protocol trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent. Condition (4) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
$\tilde{\mu}_{jk}(\bm{X})=\hat{\mu}_{jk}(\bm{X})+\frac{1}{n_{jk}(S)}\sum_{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S}\{Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i})\},$ (7)
where $n_{jk}(S)$ is the number of elements in $\{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S\}$. The corresponding limit of $\tilde{\mu}_{jk}(\bm{X})$ in Assumption 2 is $\mu_{jk}(\bm{X})+E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}$ satisfying (4) with $\delta_{jk}=\delta_{kj}=0$. Conditions (5)-(6) hold when working models are correct. Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (5)-(6) embedded. For example, (4)-(6) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2025); see S3 in the Supplement for details.
In fact, (4)-(6) always hold for the SAIPW estimator with $\bm{X}=S$, i.e., the SAIPW estimator with $\hat{\mu}_{jk}(\bm{X})=\bar{Y}_{j}(S)$, which is the sample mean of $Y_{i}$’s for $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}=S$. This is because, with $S$ as the stratum indicator, the limit of $\bar{Y}_{j}(S)$ is $E_{jk}(Y^{(j)}\mid\bm{X}=S)$, which is always a correct working model, and consequently conditions (4)-(6) hold. Therefore, the SAIPW estimator with $\bm{X}=S$ is asymptotically more efficient than the SIPW estimator according to Corollary 3 and it is used in later corollaries to bridge the comparison between IPW-type and PS-type estimators.
Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in some trivial scenarios.
Corollary 4 (Comparison of PS and APS).
Under Assumptions 1-2 and (4)-(6), the APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,apw}$ is asymptotically more efficient than the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}\!=\!E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}\Big{]}$
is positive definite, unless either one of ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ conditioned on $S$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}$ is positive semidefinite.
To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting. For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$. The following result shows the asymptotic equivalence between these two estimators.
Corollary 5 (Comparison of PS and SAIPW with $\bm{X}=S$).
Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix.
The last corollary compares APS and SAIPW with the same covariate adjustment. It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (4), the APS estimator is asymptotically equivalent to the SAIPW estimator.
Corollary 6 (Comparison of APS and SAIPW).
Under Assumptions 1-2 and condition (4), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix.
It follows from Corollaries 4-6 that, under (4)-(6), SAIPW with $\bm{X}\supset S$ is more efficient than SAIPW with $\bm{X}=S$; see Section S3 in the Supplement for a proof.
All the efficiency comparisons are summarized in Figure 5.
Figure 5: Efficiency comparisons among all robust estimators, with the corresponding corollary annotated for each comparison.
Lastly, we discuss the data used by the estimation methods as an informal way to conceptualize the efficiency comparison. When comparing treatments $j$ and $k$, all methods can robustly utilize the outcome data from arms $j$ and $k$ within ${{{\cal I}}}_{jk}$, which may achieve substantial efficiency gains compared to traditional stand-alone trials. Furthermore, covariate adjustment methods (including PS, APS, AIPW, and SAIPW) can leverage additional strata and covariate information from all individuals within ${{{\cal I}}}_{jk}$, even those assigned to arms other than $j$ and $k$. This can lead to further efficiency gains, especially in platform trials with many arms.
6 Results
As reviewed in Section 2.1 and shown in Figure 1, the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study. After completing the first study, participants on both HS and DA therapies may re-enroll in the other study contingent on eligibility and willingness to consent. The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
In our analysis, the ECE population for evaluating the effect of discontinuing HS versus continuing is the population of all individuals taking only HS or both HS and DA. Accordingly, the treatment effect of interest, $\theta_{21}-\theta_{12}$, is defined on this ECE population. A similar interpretation applies to $\theta_{31}-\theta_{13}$. Demographic and baseline characteristics for participants in the two ECE populations are presented in Tables S1-S2 of the Supplement. Note that our estimand differ slightly from those in Mayer-Hamblett et al., (2023). They analyzed the HS and DA studies separately according to their primary hypotheses of interest and included data collected after participants re-enrolled in an additional study. In contrast, we excluded data recorded after re-enrollment, as our methods currently do not account for repeated measures. We plan to address this in future research. We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
For the defined estimands, we consider the IPW, SIPW, SAIPW, PS, and APS estimators, as described in Section 4, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section S2 of the Supplement. Here, $Z$ includes baseline use of HS, DA, and enrollment windows. For SAIPW and APS, $\hat{\mu}_{jk}$ is estimated by fitting a linear regression of $Y$ on covariates using data from treatment arm $j$ in $\mathcal{I}_{jk}$. The covariates include $Z$ and several other baseline variables: two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, heterozygous, or other/unknown). As shown in Corollary 2 and simulations, when adjusting for covariates using linear working models, the AIPW estimator is numerically nearly identical to the SAIPW estimator and thus omitted. We also include the naive ANOVA as described in Section 4.2, which is subject to confounding bias. Results are shown in Figure 6.
Figure 6: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial. Naive refers to the naive ANOVA method described in Section 4.2. ANOVA* and ANCOVA* are estimates based on data from each individual sub-study. IPW, SIPW, SAIPW, PS, and APS are the correct estimators for the estimand defined on the ECE population.
For all five proposed estimators (IPW, SIPW, SAIPW, PS, and APS), all point estimates are similar, as they target the same estimand defined on the ECE population. Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 6), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA – consistent with the findings reported in Mayer-Hamblett et al., (2023). IPW and SIPW produce similar confidence intervals, likely because the true effect is near zero in this non-inferiority study. SIPW and SAIPW, as well as PS and APS, also yield similar confidence intervals, as most covariates are only weakly associated with the outcome ($R^{2}$ between 0.023 and 0.1 across three arms). This likely also explains why the naive estimator does not exhibit substantial bias in this analysis.
For comparison, we also include the ANOVA* (sample means) and ANCOVA* estimators (using linear working models to adjust for the same covariates included in SAIPW and APS separately for each arm) based solely on data from each individual sub-study. As discussed in Section 3.2, these estimators target analysis populations assigned to the HS or DA study. As a result, the interpretations of these estimators differ from the five proposed estimators. Specifically, for the HS comparison, the HS study includes 18 HS-only and 222 HS+DA participants, while the sample from the ECE population includes 18 HS-only and 427 HS+DA participants. For the DA comparison, the DA study includes 139 DA-only and 205 HS+DA participants, whereas the sample from the ECE population includes 139 DA-only and 427 HS+DA participants. More detailed demographic and baseline characteristics for participants in the HS and DA studies are presented in Tables S3 and S4. The larger population difference in the DA comparison likely explains the more noticeable discrepancy in point estimates for that result (opposite sign relative to the five proposed estimators).
Moreover, as expected, the ANOVA* and ANCOVA* estimators yield wider confidence intervals compared to the five proposed estimators. This is because, for the five proposed estimators, participants on both HS and DA assigned to treatment 1 ($n=214$) contribute to both the treatment 2 vs. 1 and treatment 3 vs. 1 comparisons. The covariate-adjusted methods (SAIPW and APS) further leverage covariate information from all participants on both HS and DA ($n=427$) in both comparisons. In contrast, ANOVA* and ANCOVA* split these participants between the HS and DA sub-studies, reducing the sample size for each comparison. As a result, the five proposed estimators achieve approximately 5%–27% variance reduction relative to the sub-study-specific estimators.
Overall, the results from the SIMPLIFY trial highlight the critical importance of clearly defining the estimand in umbrella and platform trials. Without this clarity, estimators may have interpretations that depend on the randomization ratio or trial design. An additional advantage of the five proposed estimators is the potential for efficiency gains through robust use of pooled data across sub-studies.
7 Summary and Recommendations
Master protocol trials have the potential to enhance statistical efficiency by facilitating the sharing of data across treatment evaluations. However, realizing this potential requires careful consideration of two fundamental challenges: defining meaningful estimands to reflect eligibility and concurrency, and developing robust estimation methods to accommodate the treatment assignment probabilities varying with time and individual characteristics.
In this article, we present a clear framework for constructing a clinically meaningful estimand with precise specification of the population. Our main discussion centers on comparing two treatments within the ECE population, as it preserves the integrity of randomization and avoids arbitrary dependence on randomization probabilities. This framework is adaptable to various contexts, such as comparisons across multiple treatments and evaluation of treatment effects within specific subgroups or a re-weighted population. Achieving clarity on the estimand is crucial for addressing other key issues in future statistical research, including interim analyses and the use of nonconcurrent controls. To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities. We also consider model-assisted approaches for covariate adjustment. Theoretical and empirical results are given and a detailed relative efficiency comparison is in Figure 5. We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID. Lastly, our results are applicable not only to master protocol trials but also more broadly to trials with constrained or non-uniform treatment assignment.